# CombatVLA: An Efficient Vision-Language-Action Model for Combat Tasks in 3D Action Role-Playing Games

### v1

@@ 3 Tracker and Benchmark
[…]
@@
[…]
@@ 4.1 Action-of-Thought Construction
[…]
@@ 4.2 Three-Stage Progressive Learning
[…]
@@
Firstly, to better capture the correspondence between vision and action, following He et al. [15], we introduce a contrastive loss. Specifically, given an input visual image $V$ and action-of-thought data $A$, the visual [EOS] and the final [EOS] of the LLM output serve as the local representations of $V$ and $A$, respectively. Please refer to Fig.2(d) for a quick understanding. Subsequently, based on whether the actions output by the model $A_{o}$ match the actions in the label $A_{l}$, we adjust the distance between the embeddings $\hat{v}_{EOS}$ and $\hat{a}_{EOS}$, either bringing them closer or pushing them apart.
@@ 4.3 Action Execution Framework
[…]
@@ 5.2 Main Results
Training Gathering Comprehension Reasoning Avg. Time(s)$\downarrow$
@@ 5.3 Ablation Studies
[…]

### v2

@@ 3 Tracker and Benchmark
Figure 4: Visualization of action-of-thought reasoning. Given a video or frame input, the model can quickly infer the correct action under the semantic guidance of AoT. The special token $\langle\text{TRUNC}\rangle$ speeds up efficient reasoning by truncating output.
@@
where $i_{j}=\underset{i}{\arg\min}\,(t_{f_{i}}\geq t_{a_{j}})$. It ensures that each action is aligned with the nearest future frame.
@@ 4.1 Action-of-Thought Construction
Chain-of-Thought (CoT) prompting has been proven extremely effective in enhancing the complex reasoning capabilities of LLMs and MLLMs [36, 19]. Inspired by CoT, we transform the data collected from the action tracker—which includes the set of frames $F$, the set of actions AA, and their alignment—into action-of-thought data (AoT), as illustrated in Fig. 4. Specifically, the model response is formatted in JSON, which includes [action] (such as “press space”) and [explanation] (used to describe the current state of the enemy, the physical meaning of the action, etc.). Additionally, the special token $\langle\text{TRUNC}\rangle$ represents output truncation to increase inference speed.
Question: $\langle\text{IMG}\rangle$ $\langle\text{IMG}\rangle$ $\langle\text{IMG}\rangle$ Please predict the next actions based on the frame sequence. Answer: [action] $\langle\text{TRUNC}\rangle$ [explanation] $\langle\text{EOS}\rangle$.
@@ 4.2 Three-Stage Progressive Learning
Stage3: Fine-Grained Frames-Truncated-AoT Tuning. The inference time of LLM/VLM-based models is proportional to token length due to the next token prediction requirement. Therefore, we designed a truncation strategy to mitigate the time increase associated with the introduction of AoT. As shown in Fig.4, we reorganized the AoT data by introducing a special token $\langle\text{TRUNC}\rangle$. During real-time operations, any response following $\langle\text{TRUNC}\rangle$ will be truncated. This strategy allows our model to retain the benefits provided by AoT while accelerating the inference process.
@@
Firstly, to better capture the correspondence between vision and action, following He et al. [15], we introduce a contrastive loss. Specifically, given an input visual image $V$ and action-of-thought data AA, the visual [EOS] and the final [EOS] of the LLM output serve as the local representations of $V$ and AA, respectively. Please refer to Fig.2(d) for a quick understanding. Subsequently, based on whether the actions output by the model $A_{o}$ match the actions in the label $A_{l}$, we adjust the distance between the embeddings $\hat{v}_{EOS}$ and $\hat{a}_{EOS}$, either bringing them closer or pushing them apart.
@@ 4.3 Action Execution Framework
Truncated Inference and Execution. During inference, we monitor each new output token and stop when we see the $\langle\text{TRUNC}\rangle$ token, converting prior tokens into actions. This strategy accelerates the inference speed. Subsequently, we translate the actions into Python code using the “pyautogui” library to automate mouse and keyboard controls, enabling the game character to execute combat tasks.
@@ 5.2 Main Results
[…]
@@ 5.3 Ablation Studies
Ablation of Adaptive Loss. As progressive learning proceeds, the model’s performance in gathering and reasoning gradually improves. Due to the introduction of the $\langle\text{TRUNC}\rangle$ token in $Stage3$, which places the action’s explanation after the action and truncates it, the model cannot access the semantic information when generating actions, thus somewhat impairing the model’s understanding performance. However, as shown in Tab. 5, the introduction of $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{align}}$ enhanced the model’s reasoning performance, reaching 69.71, which is 6.85 points higher than $Stage2$, with an average score increase of 2.18 points.